OpenStudy (cenaida):

Solve the system using substitution. What is the x- or y-coordinate of the solution? -2x - y = 3 3x + y = 2 A. y = 5 B. y = 7 C. y = 15/2 D. y = -13

OpenStudy (e.mccormick):

Know how to do substitution?

OpenStudy (e.mccormick):

To substitute is to put in place of. In general you solve for one of the variable, x or y, and then you put the right hand side of that solution in place of the variable on the OTHER equation: In the case of: \(ax+by=c\) \(dx+ey=f\) I can do something like this: \(dx+ey=f\) \(dx+ey-ey=f-ey\) \(dx=f-ey\) \(\dfrac{dx}{d}=\dfrac{f-ey}{d}\) \(x=\dfrac{f-ey}{d}\) Then I know what x is in terms of y. I can then replace that in the other equation: \(ax+by=c\) becomes \(a\left(\dfrac{f-ey}{d}\right)+by=c\) Then you can find y.

OpenStudy (cenaida):

is it D?

OpenStudy (e.mccormick):

Yah.